# DeepMath-103K: A Large-Scale, Challenging, Decontaminated, and Verifiable Mathematical Dataset for Advancing Reasoning

### v1

@@ 3 Construction of DeepMath-103K
4 Effectiveness of DeepMath-103K
4.1 Results on Challenging Mathematical Benchmarks
@@
Math Bench
Qwen2.5-Math-7B- Instruct 83.2 59.2 42.6 41.7 12.1 11.0
Qwen2.5-7B- Base 54.8 35.3 27.8 16.2 7.7 5.4
     Supervised Fine-Tuning
      1 R1 Solution 69.2 47.3 35.9 29.8 12.3 8.7
      3 R1 Solutions 74.1 50.0 40.2 34.1 13.8 14.0
     Zero Reinforcement Learning
      Open-Reasoner-Zero-7B 81.8 58.9 47.9 38.4 15.6 14.4
      Qwen-2.5-7B-SimpleRL-Zoo 77.0 55.8 41.0 41.2 15.6 8.7
      DeepMath-Zero-7B (Ours) 85.5 64.7 51.0 45.3 20.4 17.5
R1-Distill-Qwen-1.5B ( Instruct) 84.7 72.0 53.1 36.6 29.4 24.8
     Reinforcement Learning
      DeepScaleR-1.5B-Preview 89.4 80.3 60.9 42.2 42.3 29.6
      Still-3-1.5B-Preview 86.6 75.8 55.7 38.7 30.8 24.6
      DeepMath-1.5B (Ours) 89.0 81.6 60.1 40.6 39.8 30.8
Table 3: Results of RL and SFT using different training datasets. “DeepMath” denotes models trained using DeepMath-103K. Performance is evaluated using pass@1 (n=16) accuracy. We also list the results of Qwen2.5-Math-7B-Instruct for reference.
To empirically validate the effectiveness of DeepMath-103K for training advanced mathematical reasoning models, we conducted experiments utilizing different training paradigms. We trained models starting from two distinct initial checkpoints: a base LLM Qwen-2.5-7B-Base and a model already supervised fine-tuned for mathematical reasoning – R1-Distill-Qwen-1.5B. We evaluated performance across a suite of challenging mathematical benchmarks: MATH500 (Hendrycks et al., 2021b), AIME 2024-2025 (MAA, a), AMC 2023 (MAA, b), Minerva Math (Lewkowycz et al., 2022), and OlympiadBench (He et al., 2024). The results summarized in Table 3 clearly demonstrate the benefits of training with DeepMath-103K.
Supervised fine-tuning on DeepMath-103K significantly enhances base model performance, with multiple solutions yielding further gains.
As detailed in Section 2, each problem in DeepMath-103K includes three distinct R1-generated solutions, facilitating supervised fine-tuning (SFT). We fine-tuned Qwen-2.5-7B-Base using either the first R1 solution or all three solutions provided in DeepMath-103K. As shown in Table 3, SFT yields substantial improvements over the base model across all benchmarks. For instance, using just one R1 solution boosts MATH500 accuracy from 54.8% to 69.2%. Leveraging all three distinct solutions further improves performance (e.g., 74.1% on MATH500, 14.0% on AIME25), demonstrating the value of exposing the model to diverse problem-solving strategies.
DeepMath-103K enables state-of-the-art performance for the base model via RL-Zero.
The RL-Zero (Guo et al., 2025) trains models using online RL with binary rewards derived directly from verifiable final answers. This approach aligns perfectly with DeepMath-103K’s core design principle of providing a rigorously verified final answer for every problem (Stage 4 in Section 3), enabling direct application of rule-based rewards crucial for advancing mathematical reasoning. We trained the Qwen2.5-7B-Base model using RL-Zero on DeepMath-103K (denoted ”DeepMath-Zero”). For comparison, we included results from models trained using RL-Zero on other public RL datasets: Open-Reasoner-Zero-7B (Hu et al., 2025) and Qwen-2.5-7B-SimpleRL-Zoo (Zeng et al., 2025a) for the 7B model. Table 3 shows that DeepMath-Zero-7B consistently and significantly outperforms both the SFT models and models trained on other RL datasets across all benchmarks.
In addition, we also conducted experiments on R1-Distill-Qwen-1.5B, a model already supervised fine-tuned for mathematical reasoning. For comparison, we included results of DeepScaleR-1.5B-Preview (Luo et al., 2025b) and Still-3-1.5B-Preview (Face, 2025) for the 1.5B model. As shown in Table 3, DeepMath-Zero-1.5B achieves state-of-the-art results among the compared 1.5B models on AMC23 (81.6%) and AIME25 (30.8%), and performs competitively at the top on the remaining benchmarks.
The significant performance gains validate DeepMath-103K’s design principles for advancing mathematical reasoning.
These results strongly demonstrate DeepMath-103K’s effectiveness. The substantial improvements, particularly with RL-Zero, underscore the value of DeepMath-103K’s core characteristics. The dataset’s focus on high-difficulty problems pushes models towards more sophisticated reasoning. The availability of verifiable final answers enables effective optimization via rule-based RL. Furthermore, the rigorous decontamination process ensures that the observed gains reflect genuine reasoning improvements, not benchmark leakage. The scale and topical diversity likely contribute to the robustness observed across various benchmarks. Overall, these experiments confirm that DeepMath-103K serves as a powerful resource for training highly capable mathematical reasoning models.
4.2 Analysis of RL-Zero Using DeepMath-103K
In this section, we present a qualitative analysis to provide insights into how training on DeepMath-103K with RL-Zero enhances mathematical reasoning capabilities, focusing on DeepMath-Zero-7B.
Figure 7: Average reasoning length (number of tokens) on different benchmarks for models trained with RL-Zero using various datasets, starting from Qwen-2.5-7B-Base.
Training on DeepMath-103K often encourages models to generate substantially longer and more detailed reasoning steps, particularly on highly complex benchmarks.
We analyzed the average length (in tokens) of the solutions generated by DeepMath-Zero-7B compared to the base model and models trained on other RL datasets (Figure 7). DeepMath-Zero-7B produces markedly longer reasoning chains on several benchmarks, especially the more challenging ones like AIME24 (12106 tokens vs. 4503 for Open-Reasoner-Zero and 1582 for Qwen-2.5-7B-SimpleRL-Zoo), AIME25 (7920 vs. 3177 and 1365). While solution length is not a direct measure of quality, this significant increase suggests that training on DeepMath-103K’s challenging problems prompts the model to engage in more elaborate, step-by-step derivations, potentially reflecting deeper processing required to solve complex tasks. SimpleRL-Zoo, focusing on simpler problems, leads to the shortest solutions, aligning with expectations.
Figure 8: Variance (change from initial state) of beneficial cognitive behaviors observed during RL-Zero training on DeepMath-103K.
Models trained on DeepMath-103K exhibit increased use of beneficial cognitive behaviors during RL-Zero training.
We followed the methodology of Gandhi et al. (2025); Zeng et al. (2025a) to track the emergence of four key cognitive behaviors associated with effective problem-solving during the RL-Zero training of DeepMath-Zero-7B:
  • Subgoal Setting: Breaking down complex problems into smaller, manageable steps (e.g., ”To solve this, we first need to…”).
  • Verification: Systematically checking intermediate results or reasoning steps (e.g., ”Let’s verify this result by…”).
  • Backtracking: Explicitly revising approaches upon identifying errors or dead ends (e.g., ”This approach won’t work because… Let’s try…”).
  • Enumeration: Solving problems by exhaustively considering multiple cases or possibilities.
Figure 8 shows the change in the frequency of these behaviors compared to the initial model over the course of training. We observe notable increases in all beneficial cognitive behaviors as training progresses. This suggests that optimizing on DeepMath-103K’s challenging and verifiable problems encourages the model to adopt more structured and robust problem-solving strategies, moving beyond simple pattern matching towards more deliberate reasoning processes involving planning, checking, and adapting its approach. The emergence of these behaviors aligns with the goal of developing more general and capable mathematical reasoners.
5 Related Work
LLM Complex Reasoning
LLMs use Chain-of-Thought (CoT) prompting (Wei et al., 2022) to generate step-by-step reasoning, significantly improving their effectiveness in reasoning-intensive tasks. A line of strategies tends to directly improve the LLM CoT performance on reasoning tasks through demonstrations (Zhou et al., 2022; Zhang et al., 2023; Brown et al., 2020; Wang et al., 2022), while other works contributes to the CoT structures (Yao et al., 2023; Chen et al., 2023), planning (Wang et al., 2023), and difficulty (Fu et al., 2023).
Recent advancement in inference-time scaling (Snell et al., 2024; Brown et al., 2024; Wu et al., 2024) significantly boosts the performance of LLM reasoning on complex tasks. Building on this, online reinforcement learning (RL) with correctness rewards has emerged as a promising direction for reasoning models (Guo et al., 2025). By training models to optimize for correct final answers, these methods yield state-of-the-art performance on complex reasoning tasks (Jaech et al., 2024; Guo et al., 2025; Team, 2024; xAI, 2025; Google, 2025). Motivated by this, we propose DeepMath-103K dataset of large-scale and verifiable answers, which is suitable for RL of LLM complex reasoning capabilities.
Mathematical Reasoning Datasets and Beyond
A typical paradigm to enhance the LLM reasoning capabilities is to conduct supervised fine-tuning on human-annotated reasoning data. Those mathematical reasoning data of question-answer pairs are mainly mined from the web corpus (LI et al., 2024; Zeng et al., 2024; Li et al., 2024; Yu et al., 2024; Yue et al., 2024). Correspondingly, with improved reasoning capabilities of LLMs, the evaluation benchmark datasets focus on assessing LLMs on problem-solving accuracy, step-by-step reasoning, and robustness on more challenging and difficulty levels beyond the human experts (Hendrycks et al., 2021b; Lewkowycz et al., 2022; He et al., 2024; Gao et al., 2024; Fang et al., 2024; Zhong et al., 2023; Arora et al., 2023; Hendrycks et al., 2021a; Wei et al., 2023; Huang et al., 2024; Cobbe et al., 2021a; Rein et al., 2024; Cobbe et al., 2021b; MAA, a; b; Glazer et al., 2024).
In the RL-Zero era, the scarcity of challenging, diverse, and verifiable reasoning questions impedes reinforcement learning-based zero-shot LLM reasoning. To address this, we introduce DeepMath-103K, a novel dataset of carefully curated questions spanning varied difficulty levels and topics. Additionally, DeepMath-103K  is designed to be contamination-free with respect to evaluation benchmarks, mitigating concerns about test set leakage.
RL-Zero Reinforcement Learning
RL-Zero (Guo et al., 2025) is a streamlined framework for developing reinforcement learning (RL) capabilities in base LLMs from scratch. It elucidates LLM reasoning by employing RL principles, using interactions with the environment, i.e., the correctness of reasoning on question examples, to bolster exploration and enhance reasoning performance. Recent progress in large language model (LLM) reasoning has been profoundly shaped by reinforcement learning (RL) techniques (Jaech et al., 2024; Guo et al., 2025; Team, 2024; xAI, 2025; Google, 2025).
Despite the successes of RL-Zero in advancing LLM reasoning, models tuned within this framework face challenges in reasoning efficiency. As observed by Chen et al. (2024) and Wang et al. (2025), these models tend to overthink simple questions and generate redundant CoTs through excessive verification, while underthinking complex ones leading to inconsistent performance across diverse problem difficulties. Based on the above analysis, recent studies focus on developing efficient RL-zero algorithms to enhance the LLMs’ efficient reasoning capabilities (Chen et al., 2024; Team et al., 2025; Wang et al., 2025; Arora & Zanette, 2025; Yeo et al., 2025; Luo et al., 2025a). A core principle in these efforts is calibrating the LLM’s response length to the actual difficulty of the task, which motivates us to construct DeepMath-103K dataset with difficulty annotation.
6 Conclusion
In this work, we addressed the critical bottleneck hindering progress in AI mathematical reasoning, particularly for reinforcement learning approaches: the lack of large-scale, sufficiently challenging, verifiable, and clean training data. We introduced DeepMath-103K, a novel dataset meticulously constructed to overcome these limitations. DeepMath-103K offers a substantial volume ( 103K problems) of primarily high-difficulty mathematical problems (Levels 5-9), each equipped with a verifiable final answer essential for rule-based RL, and further enriched with three distinct R1-generated solutions to support diverse training methodologies. Our rigorous curation process, featuring comprehensive decontamination against standard benchmarks, ensures the dataset’s integrity for reliable model evaluation. We demonstrated the practical value of DeepMath-103K, showing that models trained on it exhibit significant performance gains on demanding mathematical reasoning tasks. By publicly releasing DeepMath-103K, we provide a vital resource to the research community, aimed at accelerating the development of more powerful and robust AI systems capable of tackling complex mathematical challenges.

### v2

@@ 3 Construction of DeepMath-103K
4 DeepMath Series Models
This section presents a comprehensive evaluation of the mathematical and general reasoning capabilities of our DeepMath series of models, which were trained on DeepMath-103K.
4.1 Setup
Training Paradigms
We employed two distinct RL training paradigms:
  • Zero RL: This paradigm involves training LLMs from their base (non-instruction-tuned) version using RL (Guo et al., 2025). We used group relative policy optimization (GRPO) (Shao et al., 2024) with fixes from Yu et al. (2025), and trained Qwen-2.5-(Math)-7B with a rule-based reward (+1 for correct final answer, -1 otherwise). Detailed settings are available in Footnote 3, and SFT results are presented in Appendix B.
  • RL: We also performed RL on instruction-tuned models that already possessing math reasoning ability. We explored this using R1-Distill-Qwen-1.5B (Guo et al., 2025) and OpenMath-Nemotron-1.5B (Moshkov et al., 2025).
Evaluation
Following Zeng et al. (2025a; b), we assessed the mathematical performance of our models on: MATH-500 (Hendrycks et al., 2021b), AMC 2023 (MAA, b), OlympiadBench (He et al., 2024), Minerva Math (Lewkowycz et al., 2022), AIME 2024-2025 (MAA, a), and the English subset of PolyMath (Wang et al., 2025a). To investigate the generalization of reasoning abilities beyond mathematics, we used the GPQA-Diamond benchmark, which covers biology, physics and chemistry (Rein et al., 2024). For all evaluations, we adopted pass@1 accuracy (averaged over 16 samples) as the metric, and fixed the decoding parameters to temperature=0.6, top_p=0.95, and max_tokens=32K. To ensure a fair comparison and eliminate variance caused by the evaluation script, we re-evaluated the performance of all baseline models under our evaluation settings.
4.2 Mathematical Reasoning Results
Table 3: Math reasoning performance. “DeepMath” denotes models trained on DeepMath-103K.
Model MATH AMC Olympiad Minerva AIME AIME Poly
500 23 Bench Math 24 25 Math
Proprietary Models
o1-mini 63.6
o3-mini (low effort) 60.0
Zero RL from Base Model
Qwen-2.5-7B (Team, 2024) 54.8 35.3 27.8 16.2 7.7 5.4 28.1
[…]
Qwen-2.5-Math-7B (Team, 2024) 46.9 31.9 15.8 15.5 11.2 4.4 22.7
[…]
RL from Instruct Models
R1-Distill-Qwen-1.5B (Guo et al., 2025) 84.7 72.0 53.1 36.6 29.4 24.8 39.9
[…]
OpenMath-Nemotron-1.5B (Moshkov et al., 2025) 91.8 90.5 70.3 26.3 61.3 50.6 56.8
[…]
The results presented in Table 3 collectively demonstrate the effectiveness of DeepMath-103K as a valuable resource for advancing the state-of-the-art in mathematical reasoning:
Zero RL Training on Base Model
DeepMath-Zero-7B and DeepMath-Zero-Math-7B, trained from the base Qwen-2.5-7B and Qwen-2.5-Math-7B models, demonstrate significant performance gains and achieve new SOTA results on all evaluated benchmarks. These results highlight the effectiveness of DeepMath-103K in enabling the training of powerful mathematical reasoners from scratch.
RL Training on Instruction-tuned Models
Fine-tuning instruction-tuned models with RLVR on DeepMath-103K also yields notable performance enhancements. DeepMath-1.5B, initialized from R1-Distill-Qwen-1.5B, achieves strong performance, particularly on AMC23 (82.3%) and OlympiadBench (61.8%). Similarly, DeepMath-Omn-1.5B, starting from OpenMath-Nemotron-1.5B, attains new SOTA results among 1.5B-scale models on all evaluated benchmarks, and even surpasses o1-mini and o3-mini (low effort). The consistent improvements observed across different instruction-tuned baselines further validate the utility of DeepMath-103K in boosting strong models.
4.3 Generalizable Reasoning beyond Mathematics
Table 4 presents the reasoning performance of DeepMath models on the GPQA-Diamond (Rein et al., 2024), which covers biology, physics, and chemistry. DeepMath series models achieve superior performance compared to other baseline, demonstrating a remarkable capacity to generalize their reasoning abilities acquired from mathematics to broader domains. We attribute this generalization to the data diversity and rigorous curation of DeepMath-103K. By sourcing less structured but more diverse data like Math StackExchange, DeepMath-103K yields a dataset with unique and diverse problems. Furthermore, the rigorous curation pipeline ensures both the challenge and the integrity of the data. This exposure to a wider variety of problem-solving scenarios and reasoning styles likely equips our models with more robust and transferable reasoning skills.
Table 4: Performance on the GPQA-Diamond benchmark.
Model Biology Physics Chemistry Overall
Zero RL from Base Model
Qwen-2.5-7B 33.6 27.8 21.4 25.3
[…]
Qwen-2.5-Math-7B 32.2 26.0 21.1 24.3
[…]
RL from Instruct Models
R1-Distill-Qwen-1.5B 13.5 36.2 4.4 19.1
[…]
OpenMath-Nemotron-1.5B 12.8 23.5 18.9 20.3
[…]
4.4 Analysis of Zero RL Using DeepMath-103K
Figure 9 presents an analysis of the characteristics observed during the zero RL training of DeepMath-Zero-7B. Specifically, Figure 9a illustrates the trend of rollout response length throughout the training process, while Figure 9b tracks the emergence of four cognitive behaviors following the method outlined by Gandhi et al. (2025) and Zeng et al. (2025a). The increasing trends in both response length and the manifestation of cognitive behaviors suggest a reproduction of the “aha moment” phenomenon observed in DeepSeek-R1 (Guo et al., 2025), which leads to a long reasoning model. Furthermore, Figure 9c shows the average response lengths of different models on the evaluated benchmarks. The notably longer response lengths exhibited by DeepMath-Zero-7B indicate that DeepMath-103K serves as a valuable resource for research on long reasoning models, particularly concerning phenomena such as over- and under-thinking (Chen et al., 2024; Wang et al., 2025b).
(a) Response length (training)
(b) Change of cognitive behaviors.
(c) Response length (test)
Figure 9: (a) DeepMath-103K is challenging compared to existing datasets. (b) Results of zero RL and RL using DeepMath-103K on AIME25.
5 Related Work
Datasets for advancing mathematical reasoning of LLM falls into three main lines corresponding to the three stages of LLM post-training: continue pre-training (CPT), SFT and RL. CPT aims to inject fundamental mathematical knowledge into LLMs with representative works like OpenWebMath (Paster et al., 2023), MathPile (Wang et al., 2024), InfiMM-Web-Math (Han et al., 2024), FineMath (Allal et al., 2025), and MegaMath (Zhou et al., 2025). SFT has been a foundational approach, utilizing datasets like MATH (Hendrycks et al., 2021b) and GSM8K (Cobbe et al., 2021) which provide problems with step-by-step solutions to teach models reasoning patterns. Subsequent efforts have focused on creating larger, harder and more diverse SFT datasets, such as MetaMathQA (Yu et al., 2024), OpenMathInstruct (Toshniwal et al., 2024b; a), NuminaMath-CoT (LI et al., 2024), MMIQC (Liu et al., 2024), dart-math-hard (Tong et al., 2024), and OpenMathReasoning (Moshkov et al., 2025). Recent progress in RLVR catalyzes datasets with verifiable reward, such as Open-R1 (Face, 2025), ORZ-129K (Hu et al., 2025), DSR-Preview (Luo et al., 2025), Open-RS (Dang & Ngo, 2025), DAPO-17K (Yu et al., 2025), and BigMath (Albalak et al., 2025). DeepMath-103K distinguishes itself by a unique blend of high difficulty, rigorous decontamination, and verifiable answers.
6 Conclusion
In this work, we introduce DeepMath-103K, a large-scale mathematical dataset specifically designed to advance the reasoning capabilities of LLMs through RLVR. DeepMath-103K distinguishes itself through its high concentration of challenging problems, rigorous decontamination against a wide range of benchmarks, and the inclusion of verifiable final answers and multiple diverse solutions for each problem. Our data curation pipeline leverages the richness of less structured mathematical forums, resulting in a dataset with significant novelty and diversity compared to existing resources. Our experiments demonstrate the substantial impact of DeepMath-103K. Models trained on this dataset, the DeepMath series, achieve new SOTA results on many mathematical benchmarks and exhibit remarkable generalization to domains beyond mathematics. By releasing the DeepMath-103K dataset, along with our code and model weights, we aim to provide a robust platform for the community to further explore and push the boundaries of advanced reasoning.
Appendix A Contamination Analysis of Existing Datasets
Figure 10: Number of contaminated samples in various datasets when compared against the MATH500 benchmark.
We performed a contamination analysis of several existing datasets, including ORZ-129K (Hu et al., 2025), DSR-Preview (Luo et al., 2025), DAPO-17K (Yu et al., 2025), Open-RS (Bansal et al., 2025), Open-R1 (Face, 2025), and DeepMath-103K. Our analysis focused on detecting potential contamination from the MATH500 (Hendrycks et al., 2021b), a commonly used benchmark. We employed a string-based comparison method, specifically identifying cases where the normalized indel similarity between a problem in the analyzed dataset and a problem in MATH500 exceeded 90%. This approach is notably more lenient than the rigorous semantic decontamination procedure used in the construction of DeepMath-103K (§ 3). However, the numbers of contaminated samples shown in Figure 10 reveal that most of the analyzed datasets exhibit some degree of contamination, with the exception of DeepMath-103K.
Appendix B SFT Results
As mentioned in § 2, each problem in DeepMath-103K includes three distinct R1-generated solutions, facilitating SFT. We fine-tuned Qwen-2.5-7B using either the first R1 solution or all three solutions. Table 5 shows that SFT on DeepMath-103K with one solution each problem also significantly enhances base model performance across all benchmarks, with multiple solutions yielding further gains. However, SFT still lags behind RL.
Table 5: Math reasoning performance after fine-tuning Qwen-2.5-7B via SFT. We also add DeepMath-Zero-7B as an RL counterpart for reference.
@@ Appendix B SFT Results
Bench Math
Qwen-2.5-7B 54.8 35.3 27.8 16.2 7.7 5.4
[…]
Appendix C Training Details
We use verl as the training framework [^3] https://github.com/volcengine/verl. Configurations for training DeepMath series models are listed in Table 6.
Table 6: Configurations for training DeepMath series models.
Config DeepMath-Zero-7B DeepMath-Zero-Math-7B DeepMath-1.5B DeepMath-Omn-1.5B
lr 1e-6 1e-6 1e-6 1e-6
kl_coef 0.0 0.0 1e-3 1e-3
max_prompt_length 2K 1K 2K 2K
max_response_length 10K 3K 24K 24K
train_batch_size 512 512 128 128
ppo_mini_batch_size 32 32 64 64
clip_ratio_low 0.20 0.20 0.20 0.20
clip_ratio_high 0.28 0.28 0.27 0.27
temperature 1.0 1.0 0.6 0.6
rollout.n 16 16 16 18
overlong_buffer.len 2K 512 4K 4K
total_training_steps 500 500 1800 700
Appendix D Licenses for Existing Assets
Table 7: Licenses for existing assets.
Type Asset License
Data MMIQC (Liu et al., 2024) Apache 2.0
WebInstSub (Yue et al., 2024) Apache 2.0
NuminaMath-CoT (LI et al., 2024) Apache 2.0
Code verl (Sheng et al., 2024) Apache 2.0
NeMo-Skills (Toshniwal et al., 2024b) Apache 2.0
Model Qwen-2.5-7B Team (2024) Apache 2.0
R1-Distill-Qwen-1.5B (Guo et al., 2025) MIT
OpenMath-Nemotron-1.5B (Moshkov et al., 2025) CC BY 4.0
Appendix E Limitations and Broader Impacts
While DeepMath-103K advances mathematical AI, its difficulty assessment relies on LLM evaluations, potentially introducing bias. Topical diversity may not be perfectly balanced, and the dataset’s creation was computationally intensive. Our manual analysis reveals judgment and multiple-choice questions whose answers might be matched successfully via random guess. However, DeepMath-103K’s public release can lower the barrier for RL reasoning research, accelerate progress on challenging problems, improve benchmark reliability, and foster more generalizable AI.